OpenStudy (anonymous):

could someone work out y'' for y= ln x/x^2

OpenStudy (anonymous):

here is the working

OpenStudy (anonymous):

\[y=\frac{\ln x }{x^2} \rightarrow y'=\frac{x^2 (\ln x)'-\ln x (x^2)'}{(x^2)^2}=\frac{x^2/x-2x \ln x}{x^4}\]\[=\frac{x-2x \ln x}{x^4}=\frac{x-2\ln x}{x^3}\]

OpenStudy (anonymous):

i moved the x^2 to the top so i could use the product rule

OpenStudy (anonymous):

\[y''= \left( \frac{x-2 \ln x}{x^3} \right)'=\frac{x^3(x-2 \ln x)'-(x-2\ln x)(x^3)'}{(x^3)^2}\]\[=\frac{x^3(1-2/x)-(x-2\ln x)(3x^2)}{(x^3)^2}\]\[=\frac{x^3-2x^2-3x^3+6x^2 \ln x}{x^6}=\frac{x-2-3x+6\ln x}{x^4}\]\[=\frac{6\ln x-2(x+1)}{x^4}\]

OpenStudy (anonymous):

ballards, your way is equivalent, and actually, my preferred method! I thought you might have needed it using the quotient rule.

OpenStudy (anonymous):

our teacher said use whatever way we can get it. could you work it out that way so i can check my work?

OpenStudy (anonymous):

I'm rushing through stuff. Hope it's what you've got.

OpenStudy (anonymous):

here is what i got

OpenStudy (anonymous):

the second line is the first derivative

OpenStudy (anonymous):

Your result seems fine.

OpenStudy (anonymous):

You just need to simplify now your second derivative now.

OpenStudy (anonymous):

The result in my attachment is what you should end up with. Ignore what I typed up on this thing earlier. Working on this thing is impossible.

OpenStudy (anonymous):

so could you show me how to do that. i am stuck

OpenStudy (anonymous):

It's kind of difficult to explain much further online :( Are you having a problem with the working in my attachment, or in simplifying the algebra in your own answer?

OpenStudy (anonymous):

SIMPLIFYING

OpenStudy (anonymous):

I would first multiply each of your products to get four little expressions, and then collect like terms. I started typing but it's going to take forever. I will do it on paper and scan.

OpenStudy (anonymous):

okay thank you so much!

OpenStudy (anonymous):